A company’s total sales (in millions of dollars) ? months from now are given by:?(?) = 0.015?4 + 0.4?3 + 3.4?2 + 10? − 3

Use the results from (b) to estimate the total sales 5 months from now.

Find ?′(?).

Find ?(4) and ?′(4) (round to two decimal places). Write a verbal interpretation of these results.
Answer #1

Sit ) = 0.01574 +01473 +3,4+? +106-3 775 515) = 0.015454+0.4453+3:485271045-3 = 1 91,3757 million dollag 1 S(t) = 0.015X4+3+